OpenStudy (anonymous):

x=2y+3z if your solving for y???

mathslover (mathslover):

Subtract 3z both sides

mathslover (mathslover):

\[\large{x-\color{red}{3z}=\color{blue}{2y+3z}-\color{red}{3z}}\]

mathslover (mathslover):

\[\large{\color{blue}{x-3z}=2y}\] \[\large{\frac{x-3z}{2}=\frac{2y}{2}}\]